A charge Q is taken, how should it divided in tow parts and kept certain distance constant so that force will be maximum

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
From the Coulomb's Law, k and r are constant in this problem. That leaves us to the product q1*q2 which has to be maximum.
Since q1 + q2 = Q = const, q2 = Q - q1.
So q1*(Q - q1) = max.
Let the derivative equal to zero and find the solution for q1. It is easy to probe, for this case, that the critical point is the maximum for this function.
The answer is q1 = q2 = Q/2.

Regards
Arun (askIITians forum expert)